Two light bulbs are 1.0m apart.

From what distance can these light bulbs be marginally resolved by a small telescope with a 3.0cm -diameter objective lens? Assume that the lens is limited only by diffraction and ?=600nm.
Answer #1

Raleighs' criteria for angular resolution of a circular aperature of diameterD is:

? = 1.22?/D = 1.22*600x10-9 m / 3*10-2 m  = 2.44*10-5

? = d/L

L = d/? = 1.0 m /2.44*10-5 = 0.40983 *105 m

L = 40.98 Km